# TreeRPO: Tree Relative Policy Optimization

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable reasoning capabilities through Reinforcement Learning with Verifiable Rewards (RLVR) methods. However, a key limitation of existing approaches is that rewards defined at the full trajectory level provide insufficient guidance for optimizing the intermediate steps of a reasoning process. To address this, we introduce TreeRPO, a novel method that estimates the mathematical expectations of rewards at various reasoning steps using tree sampling. Unlike prior methods that rely on a separate step reward model, TreeRPO directly estimates these rewards through this sampling process. Building on the group-relative reward training mechanism of GRPO, TreeRPO innovatively computes rewards based on step-level groups generated during tree sampling. This advancement allows TreeRPO to produce fine-grained and dense reward signals, significantly enhancing the learning process and overall performance of LLMs. Experimental results demonstrate that our TreeRPO algorithm substantially improves the average Pass@1 accuracy of Qwen-2.5-Math on test benchmarks, increasing it from 19.0% to 35.5%. Furthermore, TreeRPO significantly outperforms GRPO by 2.9% in performance while simultaneously reducing the average response length by 18.1%, showcasing its effectiveness and efficiency. Our code will be available at https://github.com/yangzhch6/TreeRPO.
@@ 1 Introduction
Recent advancements in test-time scaling with reinforcement learning methods bring milestone progress to Large Language Models (LLMs). Reasoning models such as OpenAI O1 (OpenAI, 2024), DeepSeek R1 (Guo et al., 2025), and QwQ (Qwen, 2024) have demonstrated significantly superior performance in complex reasoning tasks. Reinforcement Learning with Verifiable Rewards (RLVR) plays a pivotal role in this progress, which enhances the model’s performance by continuously exploring reasoning paths toward correct answers on verifiable problems.
@@
  • Through extensive experimentation, TreeRPO was found to significantly increase Qwen-2.5-Math’s Pass@1 accuracy on various benchmarks from 19.0% to 35.5%, including a 2.9% lead over GRPO.
@@ 2.2 Reinforcement Learning with LLMs
Reinforcement Learning from Human Feedback (RLHF; Ouyang et al. 2022) has been widely used in LLM alignments. Direct Preference Optimization (DPO; Rafailov et al. 2023) is further proposed to simplify the training pipeline of RLHF, which directly uses pair-wise preference data for model optimization. Recent studies (OpenAI, 2024; Guo et al., 2025; XAI, 2024; DeepMind, 2024; Qwen, 2024; Team et al., 2025) have shown that reinforcement learning can significantly improve the reasoning ability of models. This type of work can roughly be divided into two categories:
@@
2) Reward model-free: DPO is one of these methods, but it requires the elaborate construction of pairwise data for training. Step-DPO (Lai et al., 2024) constructs a pipeline to generate pair-wise step-level data and surpasses the performance of DPO. The other line of research (Shao et al., 2024; Hu et al., 2025; Zeng et al., 2025; Luo et al., 2025b) has shown that verification functions are effective in improving the reasoning capabilities of LLMs. They avoid the need for reward models, offering a simple yet effective approach. The typical methods are GRPO (Shao et al., 2024) and its variants DAPO (Yu et al., 2025) and VAPO (Yue et al., 2025). However, rule-based reward is similar to ORM, providing trajectory-level reward signals rather than fine-grained process reward signals. Unlike existing efforts, TreeRPO achieves fine-grained, dense reward signals without relying on a separate reward model. TreeRPO can offer a more scalable solution for obtaining dense reward signals in complex reasoning tasks.
@@ 3 TreeRPO: Methodology
Figure 2: The sampling process of our TreeRPO. TreeRPO starts from the question, sampling $N$ nodes at each step until generation is completed or the maximum depth limit $D$ is reached. Then, a verifiable reward function is used to evaluate all leaf nodes and then back-propagates the rewards to their parent nodes, thereby obtaining intermediate step rewards, which achieves process reward signaling. We traverse each node and aggregate all children steps of a node into a group to compute advantages, which are finally formatted into the training batch.
@@ 3.3 Advantage Computation
However, when applied to continuous rewards, this approach introduces significant bias. For instance, two reward sequences, $\mathbf{R}_{1}=[0,0,1,1]$ and $\mathbf{R}_{2}=[0.49,0.49,0.51,0.51]$, produce identical normalized advantages $[-1,-1,1,1]$, despite their distinct reward distributions. While $\mathbf{R}_{1}$ exhibits a clear bimodal separation, $\mathbf{R}_{2}$ contains only minor variations (a maximal difference of $0.02$). This indicates that standard normalization fails to properly scale advantages for continuous rewards, leading to misleading policy updates.
@@ 4 Experiments
4.1 Main Results
@@
We conduct TreeRPO and GRPO on Qwen2.5-Math-1.5b model with the training split of the MATH dataset, and conduct the evaluation on four selected benchmarks: Math-500, MinervaMath, OlympiadBench, and AIME. As shown in Figure 3, our TreeRPO outperform GRPO on all of the tested benchmarks. After training 360 steps, our TreeRPO outperforms GRPO by 2.7% on MATH-500, 3.5% on MinervaMath, 2.4% on OlympiadBench, and 3.0% on AIME, respectively. As illustrated in Figure 1, TreeRPO outperforms the overall performance of GRPO by 2.9%. In conclusion, TreeRPO has demonstrated consistent superiority on multiple benchmarks.
@@
We conduct TreeRPO and GRPO on the Qwen2.5-Math-1.5b model with the training split of the MATH dataset, and compute the average response length on four selected benchmarks: Math-500, MinervaMath, OlympiadBench, and AIME. As illustrated in Figure 4, compared to GRPO, our TreeRPO achieves a 17.1% reduction in token usage on MATH, 22.3% on MinervaMath, 18.0% on OlympiadBench, and 15.3% on AIME. On average, TreeRPO demonstrates a 18.1% decrease in token usage across the four benchmarks compared to GRPO, showcasing its superior efficiency.
We show the response case of a simple question in Figure 5. It can be seen that in this simple case, TreeRPO’s response is more concise
@@
In the experiments, we conduct experimental analyses using different batch sizes, and the results are shown in Figure 6. For GRPO and TreeRPO, the batch size $bsz=16/128$ has very little influence on the final performance. Our TreeRPO significantly outperforms GRPO in both two Settings.
In addition, there are 3 hyperparameters to control the generation procedure of tree sampling:
  • Branching Factor $N$: At each decoding step, the model samples N candidates.
  • Maximum Depth Limit $D$: The tree expansion terminates when any path reaches the maximum depth.
  • Maximum Step Length $L_{step}$: Each step produces at most $L_{step}$ tokens per branch.
We will conduct more detailed experimental analyses on these 3 hyperparameters in the future.
@@ 5 Conclusion
6 Future Work and Limitations

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable reasoning capabilities through Reinforcement Learning with Verifiable Rewards (RLVR) methods. However, a key limitation of existing approaches is that rewards defined at the full trajectory level provide insufficient guidance for optimizing the intermediate steps of a reasoning process. To address this, we introduce TreeRPO, a novel method that estimates the mathematical expectations of rewards at various reasoning steps using tree sampling. Unlike prior methods that rely on a separate step reward model, TreeRPO directly estimates these rewards through this sampling process. Building on the group-relative reward training mechanism of GRPO, TreeRPO innovatively computes rewards based on step-level groups generated during tree sampling. This advancement allows TreeRPO to produce fine-grained and dense reward signals, significantly enhancing the learning process and overall performance of LLMs. Experimental results demonstrate that our TreeRPO algorithm substantially improves the average Pass@1 accuracy of Qwen-2.5-Math on test benchmarks, increasing it from 19.0% to 35.5%. Furthermore, TreeRPO significantly outperforms GRPO by 2.9% in performance while simultaneously reducing the average response length by 18.1%, showcasing its effectiveness and efficiency.
@@ 1 Introduction
Recent advancements in test-time scaling with reinforcement learning methods bring milestone progress to Large Language Models (LLMs). Reasoning models such as OpenAI O1 (OpenAI, 2024), DeepSeek R1 (Guo et al., 2025a), and QwQ (Qwen, 2024) have demonstrated significantly superior performance in complex reasoning tasks. Reinforcement Learning with Verifiable Rewards (RLVR) plays a pivotal role in this progress, which enhances the model’s performance by continuously exploring reasoning paths toward correct answers on verifiable problems.
@@
  • Through extensive experimentation, TreeRPO was found to significantly increase Qwen-2.5-Math-1.5B’s Pass@1 accuracy on various benchmarks from 19.0% to 35.5%, including a 2.9% lead over GRPO.
@@ 2.2 Reinforcement Learning with LLMs
Reinforcement Learning from Human Feedback (RLHF; Ouyang et al. 2022) has been widely used in LLM alignments. Direct Preference Optimization (DPO; Rafailov et al. 2023) is further proposed to simplify the training pipeline of RLHF, which directly uses pair-wise preference data for model optimization. Recent studies (OpenAI, 2024; Guo et al., 2025a; XAI, 2024; DeepMind, 2024; Qwen, 2024; Team et al., 2025) have shown that reinforcement learning can significantly improve the reasoning ability of models. This type of work can roughly be divided into two categories:
@@
2) Reward model-free: DPO is one of these methods, but it requires the elaborate construction of pairwise data for training. Step-DPO (Lai et al., 2024) constructs a pipeline to generate pair-wise step-level data and surpasses the performance of DPO. The other line of research (Shao et al., 2024; Hu et al., 2025; Zeng et al., 2025; Luo et al., 2025b) has shown that verification functions are effective in improving the reasoning capabilities of LLMs. They avoid the need for reward models, offering a simple yet effective approach. The typical methods are GRPO (Shao et al., 2024) and its variants DAPO (Yu et al., 2025) and VAPO (Yue et al., 2025). However, rule-based reward is similar to ORM, providing trajectory-level reward signals rather than fine-grained process reward signals. VinePPO (Kazemnejad et al., 2025) conduct value estimation with Monte Carlo Tree Search. However, the auxiliary rollouts are not used in policy gradient updates, while our TreeRPO utilize all rollouts of the whole tree. Concurrently, SPO (Guo et al., 2025b) also leverages segment-level advantage estimation at an intermediate granularity, achieving a better balance by offering more precise credit assignment. In general, unlike existing efforts, TreeRPO achieves fine-grained, dense reward signals without relying on a separate reward model. TreeRPO can offer a more scalable solution for obtaining dense reward signals in complex reasoning tasks.
@@ 3 TreeRPO: Methodology
Figure 2: The sampling process of our TreeRPO. TreeRPO starts from the question, sampling NN nodes at each step until generation is completed or the maximum depth limit $D$ is reached. Then, a verifiable reward function is used to evaluate all leaf nodes and then back-propagates the rewards to their parent nodes, thereby obtaining intermediate step rewards, which achieves process reward signaling. We traverse each node and aggregate all children steps of a node into a group to compute advantages, which are finally formatted into the training batch.
@@ 3.3 Advantage Computation
However, when applied to continuous rewards, this approach introduces significant bias. For instance, two reward sequences, $\mathbf{R}_{1}=[0,0,1,1]$ and $\mathbf{R}_{2}=[0.49,0.49,0.51,0.51]$, produce identical normalized advantages [1,1,1,1][-1,-1,1,1], despite their distinct reward distributions. While $\mathbf{R}_{1}$ exhibits a clear bimodal separation, $\mathbf{R}_{2}$ contains only minor variations (a maximal difference of $0.02$). This indicates that standard normalization fails to properly scale advantages for continuous rewards, leading to misleading policy updates.
@@ 4 Experiments
Table 1: Overall performance of Pass@1 (Avg@16) performance of Qwen2.5-Math series.
Method AIME24 MATH500 Olympiad Minerva Macro Accuracy
Qwen2.5-Math-1.5B as the Base Model
GRPO Baseline 13.8 67.9 28.5 20.5 32.7
TreeRPO 16.8 ($\uparrow$+3.0) 70.7 ($\uparrow$+2.8) 30.9 ($\uparrow$+2.6) 24.0 ($\uparrow$+3.5) 35.6 ($\uparrow$+2.9)
Qwen2.5-Math-7B as the Base Model
GRPO Baseline 26.7 74.3 34.7 27.1 40.7
TreeRPO 26.7 75.5 ($\uparrow$+1.2) 35.4 ($\uparrow$+0.7) 28.1 ($\uparrow$+1.0) 41.4 ($\uparrow$+0.7)
4.1 Main Results
We show the performance comparison of GRPO baseline and our TreeRPO on Qwen2.5-Math-1.5/7B in four selected benchamrk: AIME24, MATH-500, Olympiad Benchamrk, and Minerva Math. As illustrated in Table 1, for Qwen2.5-Math-1.5B, the Macro Accuracy has improved by 2.9%. Furthermore, we consider that the reason why the improvement and repetition of Qwen2.5-Math-7B is not as significant as that of Qwen2.5-Math-1.5B lies in the fact that the MATH training data for Qwen2.5-Math-7B is too simple, resulting in the improvement of the algorithm not being significantly reflected. In generall, our TreeRPO has achieved a consistency improvement compared to GRPO baseline.
@@ 4.1 Main Results
Figure 4: Performance comparison of our TreeRPO and GRPO on the four selected benchmarks: Math-500, MinervaMath, OlympiadBench, and AIME. The experiments are conducted with Qwen2.5-Math-1.5b.
@@
We conduct TreeRPO and GRPO on Qwen2.5-Math-1.5b model with the training split of the MATH dataset, and conduct the evaluation on four selected benchmarks: Math-500, MinervaMath, OlympiadBench, and AIME. As shown in Figure 3, our TreeRPO outperform GRPO on all of the tested benchmarks. We further show the dynmaic results of Qwen2.5-Math-7B in Figure 4. After training 300 steps for Qwen2.5-Math-1.5B, our TreeRPO outperforms GRPO by 2.7% on MATH-500, 3.5% on MinervaMath, 2.4% on OlympiadBench, and 3.0% on AIME, respectively. As illustrated in Figure 1, TreeRPO outperforms the overall performance of GRPO by 2.9%. In conclusion, TreeRPO has demonstrated consistent superiority on multiple benchmarks.
@@
We conduct TreeRPO and GRPO on the Qwen2.5-Math-1.5b model with the training split of the MATH dataset, and compute the average response length on four selected benchmarks: Math-500, MinervaMath, OlympiadBench, and AIME. As illustrated in Figure 5, compared to GRPO, our TreeRPO achieves a 17.1% reduction in token usage on MATH, 22.3% on MinervaMath, 18.0% on OlympiadBench, and 15.3% on AIME. On average, TreeRPO demonstrates a 18.1% decrease in token usage across the four benchmarks compared to GRPO, showcasing its superior efficiency. TreeRPO not only demonstrates an advantage in token efficiency on Qwen2.5-Math-1.5B, but also shows an efficiency advantage on Qwen2.5-Math-7B, with an average token length that is also shorter than the GRPO baseline. We show the response case of a simple question in Figure 7. It can be seen that in this simple case, TreeRPO’s response is more concise
@@
In the experiments, we conduct experimental analyses using different batch sizes, and the results are shown in Figure 6. For GRPO and TreeRPO, the batch size $bsz=16/128$ has very little influence on the final performance. Our TreeRPO significantly outperforms GRPO in both two Settings. This fully demonstrates that our TreeRPO algorithm significantly outperforms the GRPO-baseline across different hyperparameters.
@@ 5 Conclusion
6 Reproducibility Statement
To ensure the reproducibility of our research, we have meticulously assembled a comprehensive reproducibility package as part of our supplementary materials. This package is designed to enable the seamless replication of all experiments detailed in our paper. It encompasses anonymized source code that implements the proposed model and training procedures. Additionally, we have included precise configuration files and scripts that specify all hyperparameters and the training commands necessary to reproduce our results.
Appendix
Appendix A The Use of LLMs
In the preparation of this paper, large language models (LLMs), specifically DeepSeek-V3.1 and Gemini 2.5, were used solely for the purpose of polishing the writing. The LLM was employed after the core intellectual content—including the central ideas, theoretical formulations, algorithm designs, experimental setups, and result analyses—had been fully developed by the authors. The model’s assistance was limited to rephrasing sentences for improved clarity, fluency, and conciseness. All prompts provided to the LLM contained only the authors’ original text and instructions for grammatical or stylistic improvement.
Appendix B Future Work and Limitations